OpenStudy (anonymous):

are the functions f1=x and f=sinx form a linearly independent set of vectors

OpenStudy (anonymous):

on any interval?

OpenStudy (anonymous):

we can pick an interval and show that if \[ax+b\sin(x)=0\] on that interval, then \[a=0\] and \[b=0\]

OpenStudy (anonymous):

In \[\mathbb{R}^2\] I don't believe they do. For example, if you let x=0 you have the zero vector. Which I believe makes them dependent.

OpenStudy (anonymous):

of course \[ax+b\sin(x)=0\] means for all values of x

OpenStudy (anonymous):

they are linearly independent for sure

OpenStudy (anonymous):

you certainly cannot write one as a linear combination of the other. that would be a miracle.

OpenStudy (anonymous):

But like I said, if you let x=0 then you have a solution to ax+bsin(x)=0 that is NON trivial a and b. :/

OpenStudy (anonymous):

to say that \[ax+b\sin(x)=0\] on say the interval \[(-\pi,\pi)\] means \[ax+b\sin(x)=0\] for all x in that interval.

OpenStudy (anonymous):

yes but must be for all x, not for some x

OpenStudy (anonymous):

Yeah, so since x=0 makes the a and b non-trivial doesn't that one value make them dependent automatically?

OpenStudy (anonymous):

surely you can find a value of x for which \[ax+b\sin(x)=0\] but this is an identity of functions, not a value

OpenStudy (anonymous):

not the way i understand it, no. we are considering x and sin(x) as functions.

OpenStudy (anonymous):

Okay. Then obviously they are linearly independent :P

OpenStudy (anonymous):

But i'm assuming sin(x) would have to be on an interval since it is periodic?

OpenStudy (anonymous):

if they were dependent one would be a linear combination of the other yes? in other words you could write sine as a linear combination of the identity function. fat chance.

OpenStudy (anonymous):

on the interval C'(-infinity, infinity) where C' shows first deravative

OpenStudy (anonymous):

this means in the space of functions with first derivatives i believe. of course both x and sin(x) are infinitely differentiable.

OpenStudy (anonymous):

i think all you have to do is pick a number for x and show that if \[ax+b\sin(x)=0\] for that number then both a and b are zero. you can pick \[\frac{\pi}{3}\] or anything really

OpenStudy (anonymous):

how are they independent

OpenStudy (anonymous):

or you can use "wronkian" try here http://tutorial.math.lamar.edu/Classes/DE/Wronskian.aspx

OpenStudy (anonymous):

Let x=0 and that falls apart satellite. Which is what I was trying to say.

OpenStudy (anonymous):

Because that gives you a solution where a and b=/=0

OpenStudy (anonymous):

ok perhaps i am wrong. but i don't think so. for example it is well known that sine and cosine are linearly independent. but it is certainly the case that \[\cos(\frac{\pi}{4})-\sin(\frac{\pi}{4})=0\] this does not make them dependent. one is not a linear combination of the other

OpenStudy (anonymous):

the question is not "is there SOME value of x for which you can solve \[ax+b\sin(x)=0\]

OpenStudy (anonymous):

ah ok. we can in fact use wronskian. set up the 2 by 2 matrix x sin(x) 1 cos(x)

OpenStudy (anonymous):

take the determinant and get \[x\cos(x)-\sin(x)\]

OpenStudy (anonymous):

and i believe theorem says if we can find SOME value for which this determinant is non-zero, then they are independent.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Wronskian

OpenStudy (anonymous):

And doing the wronskian you have: \[\det\left[\begin{matrix}x & \sin(x) \\ 1 & \cos(x)\end{matrix}\right]=x \cos(x)-\sin(x)\] Exactly, so they are independent IFF w(x)=/=0 so you have: xcos(x)-sin(x)=0 xcos(x)=sin(x) x=tan(x) Which is only true at x=0 Which is IN THE INTERVAL. Therefore they are dependent from this point of view.

OpenStudy (anonymous):

but xcosx(x)-sinx has value of zero for x=0

OpenStudy (anonymous):

Which is what I was trying to say the whole time.

OpenStudy (anonymous):

but this is linear indepenent in howard and anton linear algebra book

OpenStudy (anonymous):

wow are we seeing the same thing but saying different things. i read it to say that if there is SOME value for which the determinant is not zero then they are independent

OpenStudy (anonymous):

not that if there is some value for which it is zero they are dependent.

OpenStudy (anonymous):

of course they are independent! one is not a linear combination of the other that is for sure. just trying for a "proof"

OpenStudy (anonymous):

Hmmm...Maybe I misunderstood. I'm trying to recall the wronskian from diffeq lol.

OpenStudy (anonymous):

the fact that they are 0 when x = 0 does not make them dependent, any more than the fact that cosine of pi/4 is the same as sine of pi/4 makes them dependent.

OpenStudy (anonymous):

Okay satellite, I concede haha.

OpenStudy (anonymous):

The Wronskian can be used to determine whether a set of differentiable functions is linearly independent on an interval. This is useful in many situations. For example, if we wish to verify that two solutions of a second-order differential equation are independent, we may use the Wronskian. Note that if the Wronskian is zero everywhere in the interval, the functions may or may not be linearly independent. A common misconception is that W = 0 everywhere implies linear dependence; the third example below shows that this is not true. However, if all of f1, ..., fn are analytic, then W = 0 everywhere implies linear dependence.

OpenStudy (anonymous):

i have confusion about x=0

OpenStudy (anonymous):

the wronskian vanishes at x = 0 but it does not vanish everywhere. that is the point.

OpenStudy (anonymous):

a set is linearly Independent if the set of solutions only contains the trivial solution x = 0, where x is the zero vector. If ANY other solution exist then they are linearly dependent

OpenStudy (anonymous):

They are independent.

OpenStudy (anonymous):

ok i got idea

OpenStudy (anonymous):

hello math girl. yes of course they are independent. but just because it is obvious doesn't mean you can't prove it

OpenStudy (anonymous):

satellite math girl is new comer

OpenStudy (anonymous):

Well, here's something that might help: if zero is a vector, k cannot be 0. If k is 0, the vector cannot be zero.

OpenStudy (anonymous):

i think this is done via wronskian or sheer obviousness. but i guess obviousness doesn't count, so wronskian does. we come up with some x for which \[x\cos(x)-\sin(x)\neq 0\] and we are finished

OpenStudy (anonymous):

i pick \[\frac{\pi}{3}\] but you can pick something else